OpenStudy (anonymous):

solve for x: square root (I don't know how to make a square root on my computer) x + 2 - 3 = 5

OpenStudy (anonymous):

I take it the 5 is not in the square root, just the x+2

OpenStudy (anonymous):

is the equation \[5=\sqrt{x-2}-3\]?

OpenStudy (anonymous):

no the equation looks like that but the =5 is after the -3

OpenStudy (anonymous):

if so then \[5=\sqrt{x+2}-3\] add 3 to both sides \[\sqrt{x-2}=8\] square both sides to remove the square root on the left \[x-2=64\] x=66

OpenStudy (anonymous):

well thats pretty much the same

OpenStudy (anonymous):

it was just easier to input into the equation generator that way